The first recorded population of a particular country was 30 million, and the population was recorded as 41 million 15 years later. The exponential growth function A=30e^k*t describes the population of this country t years later since the first recording. Use the fact that 15 years later the population increased by 11 million to find k to three decimal places.
Answer by Cromlix(4381)

Hi there,
A = 30e^kt
41 = 30e^15k
41/30 = e^15k
1.36667 = e^15k
Take natural logs of both sides
ln 1.36667 = 15k ln e
k = ln 1.36667/15
k = 0.021
Hope this helps :-)